Calculate the expected value of red marbles out of a bag containing 6 red marbles and 8 blue marbles if you are taking 4 from the bag.
Answer #1

here for hypereometric distribution parameter n=sample size =4 , m=number of red balls =6 and N =total balls =14

therefore expected value of red marbles =nm/N =4*6/14=24/14=12/7 =1.7143